Suppose you have a container filled with air at 212 oF. The volume of the container 1.00 L, the pressure of air is 1.00 atm. The molecular composition of air is 79% N2 and 21% O2 for simplification. Calculate the mass of air and moles of O2 in the container.


1
Expert's answer
2020-07-24T14:52:46-0400

n = PV/RT

  = 1×1/(0.0821×373.15)    Where, P=1 atm, T = 212 oF =100 oC = 373.15K, V=1 L

  = 0.0326 mol

XO2 = moles O2 = 0.0326×0.21 = 0.006846 (Answer)

XN2 = moles of N2 = 0.0326×0.79 = 0.025754


Mair = XO2×32 + XN2×28     

      = (0.006846×32 + 0.025754×28)g

     = 0.939 g    (Answer)

The mass of air is 0.939 g

0.006846 moles of O2 in the container.